# Diffusion Models, Image Super-Resolution And Everything: A Survey

### v2

@@ III Diffusion Models Basics
Diffusion Models (DMs) have profoundly impacted the realm of generative AI, and many approaches that fall under the umbrella term DM have emerged. At their core, these models share a fundamental principle called diffusion: generate data by gradually introducing noise into an initial, often simplistic, data sample. Next, reverse this process until it converges to the desired complex data distribution. What sets DMs apart from earlier generative models is their execution over iterative time steps, both forward and backward in time and denoted by $t$, as depicted in Figure 1. The forward and backward diffusion processes are distinguished by:
@@
Forward Diffusion $q$ degrade input data using noise iteratively, forward in time (i.e., $t$ increases).
@@
Backward Diffusion $p$ denoise the degraded data, thereby reversing the noise iteratively, backward in time (i.e., $t$ decreases).
@@
Figure 1: Principle of diffusion models. The forward diffusion adds noise iteratively (red), which translates an image from the image space to the corruption space. The backward diffusion, the iterative refinement process, reverts the process (blue) back to the image space. Shown are three different implementations of diffusion models, namely Denoising Diffusion Probabilistic Models (DDPMs), Score-based Generative Models (SGMs), and Stochastic Differential Equations (SDEs) with their respect formulation of the forward and backward diffusion.
III-A The Core Concept
Consider a dataset of LR-HR image-pairs, $\mathcal{D}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}^{N}_{i=1}$. For each time step $t$, the random variable $\mathbf{z}_{t}$ describes the current state, which constitutes an image between the image and corruption space. The time step $t$ increases incrementally during forward diffusion, whereas it propagates towards 0 during backward diffusion. In literature, there is no clear distinction between $\mathbf{z}_{t}$ in the forward diffusion and $\mathbf{z}_{t}$ in the backward diffusion. During forward diffusion, we assume $\mathbf{z}_{t}\sim q\left(\mathbf{z}_{t}\mid\mathbf{z}_{t-1}\right)$. Conversely, in the backward diffusion, we assume $\mathbf{z}_{t-1}\sim p\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{t}\right)$.
@@ III-A The Core Concept
We will denote $T$ with $0<t\leq T$ as the maximal time step for finite cases. The initial data distribution ($t=0$) is represented by $\mathbf{z}_{0}\sim q\left(\mathbf{x}\right)$, which is then slowly injected with noise (additive). Vice versa, DMs remove noise therein by running a parameterized model $p_{\theta}\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{t}\right)$ in the reverse time direction that approximates the ideal (but unattainable) denoised distribution $p\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{t}\right)$.
@@
The explicit implementation of the forward diffusion $q$ and backward diffusion $p$, approximated by $p_{\theta}$, is influenced by the specific DM in use. DMs can coarsely be divided into three types: Denoising Diffusion Probabilistic Models (DDPMs), Score-Based Generative Models (SGMs), and Stochastic Differential Equations (SDEs) [48]. Each of these types will be discussed next. The distinctive formulations of the forward and backward diffusion are also comprehensively shown in Figure 1.
III-B Denoising Diffusion Probabilistic Models (DDPMs)
@@
Denoising Diffusion Probabilistic Models (DDPMs) [4] use two Markov chains to enact the forward and backward diffusion across a finite amount of time steps.
@@
III-C Score-based Generative Models (SGMs)
Score-based Generative Models (SGMs), much like DDPMs, utilize diffusion processes but employ an alternative mathematical foundation. Instead of using probability density function $p(\mathbf{z})$ directly, Song et al. [7] propose to work with its (Stein) score function, which is defined as the gradient of the log probability density $\nabla_{\mathbf{z}}\log p(\mathbf{z})$. Mathematically, the score function preserves all information about the density function, but computationally it is easier to work with. Furthermore, the decoupling of model training from the sampling procedure grants greater flexibility in defining sampling methods and training objectives.
@@ III-C Score-based Generative Models (SGMs)
=$\stackrel{{\scriptstyle(i)}}{{=}}$ $\mathop{\mathbb{E}}_{\begin{subarray}{c}t\sim\mathcal{U}(1,T)\\ \mathbf{z}_{0}\sim q(\mathbf{z}_{0})\\ \mathbf{z}_{t}\sim q(\mathbf{z}_{t}\mid\mathbf{z}_{0})\end{subarray}}\left[\lambda(t)\sigma^{2}_{t}\|\nabla_{\mathbf{z}_{t}}\log{q(\mathbf{z}_{t}|\mathbf{z}_{0})}-s_{\theta}(\mathbf{z}_{t},t)\|^{2}\right]+c$ (28)
=$\stackrel{{\scriptstyle(ii)}}{{=}}$ $\mathop{\mathbb{E}}_{\begin{subarray}{c}t\sim\mathcal{U}(1,T)\\ \mathbf{z}_{0}\sim q(\mathbf{z}_{0})\\ \mathbf{z}_{t}\sim q(\mathbf{z}_{t}\mid\mathbf{z}_{0})\end{subarray}}\left[\lambda(t)\|-\frac{\mathbf{z}_{t}-\mathbf{z}_{0}}{\sigma_{t}}-\sigma_{t}s_{\theta}(\mathbf{z}_{t},t)\|^{2}\right]+c$ (31)
=$\stackrel{{\scriptstyle(iii)}}{{=}}$ $\mathop{\mathbb{E}}_{\begin{subarray}{c}t\sim\mathcal{U}(1,T)\\ \mathbf{z}_{0}\sim q(\mathbf{z}_{0})\\ \epsilon\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}\right)\end{subarray}}\left[\lambda(t)\|\epsilon+\sigma_{t}s_{\theta}(\mathbf{z}_{t},t)\|^{2}\right]+c$ (34)
@@
III-D Stochastic Differential Equations (SDEs)
So far, we have discussed DMs that deal with finite time steps. A generalization to infinite time steps is made by formulating these as solutions to Stochastic Differential Equations (SDEs), also known as Score SDEs [17]. In fact, we can view SGMs and DDPMs as discretizations of a continuous-time SDE. SDEs are not entirely bound to diffusion models, as they are a mathematical concept describing stochastic processes. As such, they fit perfectly to describe the processes we want to simulate in diffusion models. Like previously, data is perturbed in a general diffusion process but generalized to an infinite number of noise scales.
@@ III-D Stochastic Differential Equations (SDEs)
where f and g are the drift and diffusion functions, respectively, and $\mathbf{w}$ is the standard Wiener process (also known as Brownian motion). One benefit of this generalized formulation is that DDPMs and SGMs can be formulated uniformly in this setting. The SDE for DDPMs is given by:
@@
III-E Relation between Diffusion Models
As highlighted in the SDE section, we can describe both variations, namely SGMs and DDPMs, with SDEs. We can also showcase this close relationship by reformulating the optimization targets. For DDPMs, we saw in Equation 18 that
@@ III-E Relation between Diffusion Models
the relationship between DDPMs and SGMs becomes clear once we set $\epsilon_{\theta}(\mathbf{z}_{t},t)=-\sigma_{t}s_{\theta}(\mathbf{z}_{t},t)$. As the constant $c$ becomes irrelevant for the optimization, we can see once again that there is a mathematical connection between DDPMs and SGMs.
III-F Relation to other Image SR Generative Models
The exploration of generative models in image SR is fascinating because it offers the potential to improve the quality of upsampled images beyond interpolation significantly. They differ primarily in how they approach the task of generating HR images from LR inputs and are illustrated in Figure 2. These differences stem from the underlying architecture and training objectives. While they offer significant advantages, they come with their individual set of challenges, such as training stability and computational costs.
GAN: One prominent category of generative models is Generative Adversarial Networks (GANs) [55], which have demonstrated state-of-the-art performance in various vision-related tasks, including text-to-image synthesis [11] and image super-resolution (SR) [44]. GANs are known for their adversarial training, where a generator competes against a discriminator. While DMs do not use a discriminator in the same way, they also employ an adversarial-like training strategy, where noise is iteratively added and removed to encourage the generation of realistic data. However, approaches with GANs often suffer from non-convergence, training instability, and high computational costs. They require careful hyperparameter tuning due to the interplay between the generator and the discriminator.
@@ IV Improvements for Diffusion Models
This section introduces enhancements made to the foundational principles we have previously outlined. Two particular advancements, namely, efficient sampling and enhanced likelihood estimation, stand out due to their direct relevance and adaptability to the domain of image SR.
@@ IV-A Efficient Sampling
Efficient sampling plays a pivotal role in addressing one of the central challenges hindering the widespread adoption of DMs - the considerable computational resources required for sampling. To put it in perspective, generating samples using DMs can demand processing power equivalent to hundreds of GPU days. For instance, sampling 50,000 images of size 32x32 with a DDPM can take about 20 hours, while a GAN can achieve this in less than a minute on a Nvidia 2080 Ti GPU; the challenge escalates for larger images, as sampling 50,000 images of size 256x256 could require close to 1,000 hours on the same hardware [60]. Advanced sampling methodologies have emerged as a solution, expediting the sampling process while upholding the quality of the generated images. A key aspect of those models to achieve this is the independence of training and inference schedules. In other words, while a model may undergo training over 1000 steps, the subsequent inference phase in practice may require only a fraction. Efficient sampling strategies can be categorized into Training-based and Training-free sampling.
Training-based sampling methods speed up data generation using a trained sampler that approximates the backward diffusion process instead of a traditional numerical solver. This can be done either entirely or partially. For example, Watson et al. [61] developed a dynamic programming algorithm that identifies optimal inference paths using a fixed number of refinement steps, significantly reducing the computation required. Diffusion Sampler Search [62] offers another approach, optimizing fast samplers for pre-trained diffusion models by adjusting the Kernel Inception Distance.
@@
Knowledge distillation is also used to accelerate sampling. It involves transferring knowledge from a complex, slower sampler (the teacher model) to simpler, faster models (student models) [63, 64]. As demonstrated by Salimans et al. [65], this method progressively reduces the number of sampling steps, trading off a slight decrease in sample quality for increased speed. Similarly, Xiao et al. [66] addressed the slow sampling issue associated with the Gaussian assumption in denoising steps, which is usually only effective for small step sizes. They proposed Denoising Diffusion GANs that use conditional GANs for the denoising steps, allowing for larger step sizes and faster sampling.
Lastly, truncated diffusion is a technique that improves speed by prematurely ending the forward diffusion process [67, 68]. This early termination results in outputs that are not purely Gaussian noise, presenting computational challenges. These challenges are overcome using proxy distributions from pre-trained VAEs or GANs, which match the diffused data distribution and facilitate efficient backward diffusion.
Training-free sampling methods aim to speed up sampling by minimizing the number of discretization steps while solving the Stochastic Differential Equation (SDE) or Probability Flow Ordinary Differential Equation (ODE) [69, 70].
Denoising Diffusion Implicit Models (DDIMs) [69] introduced by Song et al. generalizes the Markovian forward diffusion of DDPMs into non-Markovian ones. This generalization allows the DDIMs to learn a Markov chain to reverse the non-Markovian forward diffusion, resulting in higher sampling speeds with minimal loss in sample quality.
In another noteworthy development, Jolicoeur-Martineau et al. [70] have devised an efficient SDE solver with adaptive step sizes for the accelerated generation of score-based models. This method has been found to generate samples more rapidly than the Euler-Maruyama method without any compromise in sample quality.
@@ IV-B Improved Likelihood
Log-likelihood improvement is directly coupled with enhancing the performance of various applications and methods, including but not limited to compression [71], semi-supervised learning [72], and image SR. Given that DMs do not directly optimize the log-likelihood, e.g., SGMs utilize a weighted combination of score-matching losses, an objective that forms an upper bound on the negative log-likelihood needs to be optimized.
Song et al. [73] proposed a method called likelihood weighting to address this need. This method minimizes the weighted combination of score matching losses for score-based DMs. A carefully selected weighting function provides an upper bound to the negative log-likelihood within the weighted score-matching objective. Upon minimization, this results in an elevation of the log-likelihood.
Kingma et al. [74] explored methods that simultaneously train the noise schedule and diffusion parameters to maximize the variational lower bound within Variational Diffusion Models.
Additionally, the Improved Denoising Diffusion Probabilistic Models (iDDPMs) proposed by Nichol and Dhariwal et al. [75] implement a cosine noise schedule. This gradually introduces noise into the input, contrasting with the linear schedules that tend to degrade the information quicker. Using the cosine noise schedule leads to better log-likelihoods and facilitates faster sampling.

### v3

@@ III Diffusion Models Basics
Diffusion Models (DMs) have profoundly impacted the realm of generative AI, and many approaches that fall under the umbrella term DM have emerged. What sets DMs apart from earlier generative models is their execution over iterative time steps, both forward and backward in time and denoted by $t$, as depicted in Figure 1. The forward and backward diffusion processes are distinguished by:
@@
Forward $q$ degrade input data using noise iteratively, forward in time (i.e., $t$ increases).
@@
Backward $p$ denoise the degraded data, thereby reversing the noise iteratively, backward in time (i.e., $t$ decreases).
@@
Figure 1: Principle of DMs. The forward diffusion adds noise iteratively (red), which translates an image from the image space to the corruption space. The backward diffusion, the iterative refinement process, reverts the process (blue) back to the image space. Shown are three different implementations of DMs, namely Denoising Diffusion Probabilistic Models (DDPMs), Score-based Generative Models (SGMs), and Stochastic Differential Equations (SDEs) with their respect formulation of the forward and backward diffusion.
@@
The time step $t$ increases during forward diffusion, whereas it propagates towards 0 during backward diffusion. Let $\mathcal{D}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}^{N}_{i=1}$ be a dataset of LR-HR image-pairs. For each time step $t$, the random variable $\mathbf{z}_{t}$ describes the current state, a state between the image and corruption space. In literature, there is no clear distinction between $\mathbf{z}_{t}$ in the forward and $\mathbf{z}_{t}$ in the backward diffusion. During forward diffusion, we assume $\mathbf{z}_{t}\sim q\left(\mathbf{z}_{t}\mid\mathbf{z}_{t-1}\right)$. Conversely, in the backward diffusion, we assume $\mathbf{z}_{t-1}\sim p\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{t}\right)$. We will denote $T$ with $0<t\leq T$ as the maximal time step for finite cases. The initial data distribution ($t=0$) is represented by $\mathbf{z}_{0}\sim q\left(\mathbf{x}\right)$, which is then slowly injected with noise (additive). Vice versa, DMs remove noise therein by running a parameterized model $p_{\theta}\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{t}\right)$ in the reverse time direction that approximates the ideal (but unattainable) denoised distribution $p\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{t}\right)$.
@@
The explicit implementation of the forward diffusion $q$ and backward diffusion $p$, approximated by $p_{\theta}$, is defined by the specific DM in use. There are three types: Two discrete forms, namely Denoising Diffusion Probabilistic Models (DDPMs) and Score-Based Generative Models (SGMs), and the continuous form by Stochastic Differential Equations (SDEs) [65]. Each of these types will be discussed next are comprehensively shown in Figure 1.
III-A Denoising Diffusion Probabilistic Models (DDPMs)
@@
Denoising Diffusion Probabilistic Models (DDPMs) [8] use two Markov chains to enact the forward and backward diffusion across a finite amount of discrete time steps.
@@
III-B Score-based Generative Models (SGMs)
Score-based Generative Models (SGMs), much like DDPMs, utilize discrete diffusion processes but employ an alternative mathematical foundation. Instead of using probability density function $p(\mathbf{z})$ directly, Song et al. [11] propose to work with its (Stein) score function, which is defined as the gradient of the log probability density $\nabla_{\mathbf{z}}\log p(\mathbf{z})$. Mathematically, the score function preserves all information about the density function, but computationally, it is easier to work with. Furthermore, the decoupling of model training from the sampling procedure grants greater flexibility in defining sampling methods and training objectives.
@@ III-B Score-based Generative Models (SGMs)
=$\stackrel{{\scriptstyle(i)}}{{=}}$ $\mathop{\mathbb{E}}_{\begin{subarray}{c}t\sim\mathcal{U}(1,T)\\ \mathbf{z}_{0}\sim q(\mathbf{z}_{0})\\ \mathbf{z}_{t}\sim q(\mathbf{z}_{t}\mid\mathbf{z}_{0})\end{subarray}}\left[\lambda(t)\sigma^{2}_{t}\|\nabla_{\mathbf{z}_{t}}\log{q(\mathbf{z}_{t}|\mathbf{z}_{0})}-s_{\theta}(\mathbf{z}_{t},t)\|^{2}\right]+c$
=$\stackrel{{\scriptstyle(ii)}}{{=}}$ $\mathop{\mathbb{E}}_{\begin{subarray}{c}t\sim\mathcal{U}(1,T)\\ \mathbf{z}_{0}\sim q(\mathbf{z}_{0})\\ \mathbf{z}_{t}\sim q(\mathbf{z}_{t}\mid\mathbf{z}_{0})\end{subarray}}\left[\lambda(t)\|-\frac{\mathbf{z}_{t}-\mathbf{z}_{0}}{\sigma_{t}}-\sigma_{t}s_{\theta}(\mathbf{z}_{t},t)\|^{2}\right]+c$
=$\stackrel{{\scriptstyle(iii)}}{{=}}$ $\mathop{\mathbb{E}}_{\begin{subarray}{c}t\sim\mathcal{U}(1,T)\\ \mathbf{z}_{0}\sim q(\mathbf{z}_{0})\\ \epsilon\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}\right)\end{subarray}}\left[\lambda(t)\|\epsilon+\sigma_{t}s_{\theta}(\mathbf{z}_{t},t)\|^{2}\right]+c$
@@
III-C Stochastic Differential Equations (SDEs)
So far, we have discussed DMs that deal with finite time steps. A generalization to infinite continuous time steps is made by formulating these as solutions to Stochastic Differential Equations (SDEs), also known as Score SDEs [10]. In fact, we can view SGMs and DDPMs as discretizations of a continuous-time SDE. SDEs are not entirely bound to DMs, as they are a mathematical concept describing stochastic processes. As such, they fit perfectly to describe the processes we want to simulate in DMs. Like previously, data is perturbed in a general diffusion process but generalized to an infinite number of noise scales.
@@ III-C Stochastic Differential Equations (SDEs)
where f and g are the drift and diffusion functions, respectively, and $\mathbf{w}$ is the standard Wiener process (also known as Brownian motion). This generalized formulation allows uniform representation of both DDPMs and SGMs. The SDE for DDPMs is given by:
@@
III-D Relation between Diffusion Models
As highlighted in the SDE section, we can describe both variations, namely SGMs, and DDPMs, with SDEs. We can also showcase this close relationship by reformulating the optimization targets. For DDPMs, we saw in Equation 18 that
@@ III-D Relation between Diffusion Models
the connection between DDPMs and SGMs becomes clear once we set $\epsilon_{\theta}(\mathbf{z}_{t},t)=-\sigma_{t}s_{\theta}(\mathbf{z}_{t},t)$. As the constant $c$ is irrelevant for the optimization, we can see once again that there is a mathematical connection between DDPMs and SGMs.
III-E Relation to other Image SR Generative Models
Generative models in image SR differ primarily in how they approach the task of generating HR images from LR inputs and are illustrated in Figure 2. These differences stem from the underlying architecture and training objectives. While they offer significant advantages, they come with a individual set of challenges, like training stability and computational costs.
GAN: One prominent category of generative models is Generative Adversarial Networks (GANs) [72], which have demonstrated state-of-the-art performance in various vision-related tasks, including text-to-image synthesis [7] and image super-resolution (SR) [52]. GANs are known for their adversarial training, where a generator competes against a discriminator. Although DMs do not employ a discriminator, they utilize a similar adversarial training strategy by iteratively adding and removing noise to enable realistic data generation. However, approaches with GANs often suffer from non-convergence, training instability, and high computational costs. They require careful hyperparameter tuning due to the interplay between the generator and the discriminator.
@@ IV Improvements for Diffusion Models
In the broader research community, there are several ways to improve DMs for image generation, as presented, for example, by Karras et al. [77]. This section, however, focuses on enhancements particularly interesting for image SR: Efficient sampling and enhanced likelihood estimation.
@@ IV-A Efficient Sampling
Efficient sampling refers to strategies that generate samples from noise more quickly, i.e., in fewer time steps, without compromising the quality of the produced image significantly. For instance, a DDPM takes about 20 hours to sample 50,000 32x32 images, in contrast to a GAN’s less than one minute on a Nvidia 2080 Ti GPU; for larger 256x256 images, this extends to nearly 1,000 hours [78]. Fortunately, the independence between training and inference schedules is often leveraged in image SR. For example, a model may undergo training with 1,000 time steps, but the subsequent inference phase may require only a fraction, i.e., 200 [15, 79]. However, the broader community of DM research has made further attempts focusing on either training-based or training-free sampling.
@@
Training-based sampling methods speed up data generation using a trained sampler that approximates the backward diffusion process instead of a traditional numerical solver. This process may be complete or partial. For example, Watson et al. [80] developed a dynamic programming algorithm that identifies optimal inference paths using a fixed number of refinement steps, significantly reducing the computation required. Diffusion Sampler Search [81] offers another approach, optimizing fast samplers for pre-trained DMs by adjusting the Kernel Inception Distance. Another technique is truncated diffusion, which improves speed by prematurely ending the forward diffusion process [82, 83]. This early termination results in outputs that are not purely Gaussian noise, presenting computational challenges. These challenges are addressed using proxy distributions from pre-trained VAEs or GANs, which match the diffused data distribution and facilitate efficient backward diffusion. Lastly, Knowledge distillation is also used to accelerate sampling. It involves transferring knowledge from a complex, slower sampler (the teacher model) to simpler, faster models (student models) [84, 85]. As demonstrated by Salimans et al. [86], this method progressively reduces the number of sampling steps, trading off a slight decrease in sample quality for increased speed. Similarly, Xiao et al. [87] addressed the slow sampling issue associated with the Gaussian assumption in denoising steps, which is usually only effective for small step sizes. They proposed Denoising Diffusion GANs that use conditional GANs for the denoising steps, allowing for larger step sizes and faster sampling. For image SR, an application for exploiting knowledge distillation can be found in AddSR [88]. Similarly, YONOS-SR [89] uses knowledge distillation, but instead of training faster samplers, they transfer different scaling task knowledge and use the training-free DDIMs for efficient sampling, which is presented in the next section.
Training-free sampling methods aim to speed up sampling by minimizing the number of discretization steps while solving the Stochastic Differential Equation (SDE) or Probability Flow Ordinary Differential Equation (ODE) [90, 91]. Denoising Diffusion Implicit Models (DDIMs) [90] introduced by Song et al. generalizes the Markovian forward diffusion of DDPMs into non-Markovian ones. This generalization allows the DDIMs to learn a Markov chain to reverse the non-Markovian forward diffusion, resulting in higher sampling speeds with minimal loss in sample quality. Jolicoeur-Martineau et al. [91] have devised an efficient SDE solver with adaptive step sizes for the accelerated generation of score-based models. This method has been found to generate samples more rapidly than the Euler-Maruyama method without compromising sample quality. Building upon DDIM and Jolicoeur-Martineau et al., the DPM-solver [92], inspired by the AnalyticalDPM [93], approximates the error prediction via Taylor expansion and thereby achieves efficient sampling by analytically resolving the linear component of the ODE solution instead of relying on generic black-box ODE solvers. This method significantly reduces the sampling steps to 10 to 20. In a later work, the authors introduced an improved version with DPM-solver++ that essentially approximates the predicted image instead of the error [94]. Lately, a more general formulation and extension of the DPM-solver++ was introduced by UniPC [95].
@@ IV-B Improved Likelihood
Log-likelihood improvement is directly coupled with enhancing the performance of various applications and methods, including but not limited to compression [96], semi-supervised learning [97], and image SR. Given that DMs do not directly optimize the log-likelihood, e.g., SGMs utilize a weighted combination of score-matching losses, an objective that forms an upper bound on the negative log-likelihood needs to be optimized. Song et al. [98] proposed a method called likelihood weighting to address this need. This method minimizes the weighted combination of score matching losses for score-based DMs. A carefully chosen weighting function sets an upper bound on the negative log-likelihood in the weighted score-matching objective. Upon minimization, this results in an elevation of the log-likelihood. Kingma et al. [99] explored methods that simultaneously train the noise schedule and diffusion parameters to maximize the variational lower bound within Variational Diffusion Models. Additionally, the Improved Denoising Diffusion Probabilistic Models (iDDPMs) proposed by Nichol and Dhariwal et al. [100] implement a cosine noise schedule. This gradually introduces noise into the input, contrasting with the linear schedules that tend to degrade the information quicker. Using the cosine noise schedule leads to better log-likelihoods and facilitates faster sampling.